RC meeting and Run Plan, Friday February 3rd

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar


RC Report

Summary

  • Swing shift, locked up Hall around 5pm. 17:30pm started bringing beam into hall. Started production data taking on polhe3 at 19:30.
  • Running production on polhe3 since then.
  • Beam was trippy during swing, but has been better during owl and day.
  • HWP is out
  • A few times after short downs beam position from nominal spiked at ~1mm, and one was even ~2mm.

Current Status

  • Currently running polhe3

Run Plan

  • Continue production running at 45uA and polHe3

Data Collection and Polarisaton Status

  • Accumulated charge uncorrected 20.3C; accumulated charge corrected 21.6C
  • Current target polarisation 49.4%

Other Topics/Status Updates

  • MCC are talking about recovery day Tue next week and accelerator tuning Thursday next week
  •  ?

Shift Schedule

Requested Accesses to the Hall

  • Looking to put SBS GEMs in DAQ during day shifts, need to fix a cable in Hall first
  • Target team for oven heater check

Special Requests

  • DC power did not access Hall during maintenance day for moeller magnet, would be good to coordinate in the future.

Notices for Shift Crew

  • Check the How To page
  • Pay attention to the beam restoration procedures in the How To
    • there is a procedure if beam has been away for 30minutes, and another procedure if machine has been down for tuning or something which may change beam profile. Communicate with MCC to figure out why beam has been down and to figure out which kind of restoration you should do.
  • SL:
    • production runs (1-1.5hr) and replays (50k replay for each run, ~two runs also with 200k replays per shift if possible)
  • TOs:
    • NMR every 4 hours
    • Note the instructions for performing an NMR if there has been any changes to magnetic fields (NMR gui needs restarted, as well as changing the holding fields).
  • Contact RC any time for any questions

Useful Links